OpenStudy (anonymous):

The measure of an interior angle of a regular polygon is 140. Find the number of sides the polygon has.

OpenStudy (anonymous):

(n-2)180=sum of the interior angles. If you just want to find one angle, then you divide (n-2)180 by the number of sides and you get = one angle of a regular polygon since you know the angle, you set the angle equal to Let the angle be represented by A. using algebra, we can say 180n-360-An=0 180n-An=360 n(180-A)=360 n=360/(180-A) Therefore, the formula to find out the number of sides of a regular polygon when one interior angle is given is: Where n=number of sides and A is the given angle.